What would be the present value of an item that has a salvage value of $25,000 at the end of five years? Assume a discount rate of 3.8% for an end-of-year factor.
1
Expert's answer
2012-12-28T04:44:21-0500
Present value, also known as present discounted value,is the value on a given date of a payment or series of payments made at other
times.
If the payments are in the future, they are discounted to reflect the time value of money and other factors such as investment risk.If they are in the past, their value is correspondingly enhanced to reflect
that those payments have been (or could have been) earning interest in the
intervening time. Present value calculations are widely used in business and
economics to provide a means to compare cash flows at different times on a meaningful "like to like" basis
PV*(1+ 0,038)^5 = 25,000
PV= 25,000/(1 + 0,038)^5 = 20,746.9013.